OpenStudy (anonymous):

Find the LCM of 4 and 20

OpenStudy (anonymous):

4

Parth (parthkohli):

\[20 = 2 \times 2 \times 5 \]\[ 4 = 2 \times 2\]Does that give a hint?

Parth (parthkohli):

@trackzz Nope, that'd make the HCF.

OpenStudy (anonymous):

woops, yeah you're right.

OpenStudy (anonymous):

What is HCF tho?

Parth (parthkohli):

HCF = GCD

OpenStudy (anonymous):

I'm loss, what is GCD?

OpenStudy (anonymous):

lost*

Parth (parthkohli):

\[ \begin{array}{l|r} 2&20,4 \\ \hline 2&10,2\\\hline 2&5,1\\\hline 5&1,1 \end{array}\]\[L.C.M = 2\times2\times2\times 5 \]

Parth (parthkohli):

@trackzz The highest number that divides both.

OpenStudy (anonymous):

The LCM is 2 right? :P

Parth (parthkohli):

No no.

Parth (parthkohli):

LCM is the lowest multiple of two numbers.

Parth (parthkohli):

The common multiples of these numbers are: 20 40 60 80 100 120 140

Parth (parthkohli):

The least number that you see here is 20. Thus, \(\text{lcm}(20,4) = 20\)

OpenStudy (anonymous):

k i need school to start now..

Parth (parthkohli):

Hahah